# Exploring the Potential of Carbon-Aware Execution for Scientific Workflows

### v1

@@ Abstract
scientific workflows, carbon-aware computing, carbon footprint, task scaling, sustainable computingI Introduction
@@ II-D Interrupted Workflow Shifting
For Chip-Seq, a potential reduction of 49% can be achieved with no overhead in the 48h window, and this improves to 64% in the 192h window with a runtime overhead of 4% – using the average signal. For Rangeland, a potential reduction of 8% can be achieved with a runtime overhead of 5% in the 48h window, improving to a reduction of 52% with a runtime overhead of 3% in the 192h window – using the marginal signal – outperforming the entire workflow shifting shifting. This exemplifies the potential for interrupted workflow shifting to offer greater reductions in carbon emissions over the same windows with a low overhead. Furthermore, more sophisticated implementations of interrupted shifting, that divide workflows into shorter task execution windows, e.g. 15min instead of 60min, and more adaptive alignment of tasks, i.e. the most energy-intensive or longest-running tasks being aligned with windows of low CI electricity, could outperform entire workflow shifting even more significantly – making better use of limited time before deadlines.

### v2

@@ Abstract
scientific workflows, carbon-aware computing, carbon footprint, task scaling, sustainable computing(17.7mm,260mm)
I Introduction
@@ II-D Interrupted Workflow Shifting
For Chip-Seq, a potential reduction of 49% can be achieved with no overhead in the 48h window, and this improves to 64% in the 192h window with a runtime overhead of 4% – using the average signal. For Rangeland, a potential reduction of 8% can be achieved with a runtime overhead of 5% in the 48h window, improving to a reduction of 52% with a runtime overhead of 3% in the 192h window – using the marginal signal – outperforming the entire workflow shifting. This exemplifies the potential for interrupted workflow shifting to offer greater reductions in carbon emissions over the same windows with a low overhead. Furthermore, more sophisticated implementations of interrupted shifting, that divide workflows into shorter task execution windows, e.g. 15min instead of 60min, and more adaptive alignment of tasks, i.e. the most energy-intensive or longest-running tasks being aligned with windows of low CI electricity, could outperform entire workflow shifting even more significantly – making better use of limited time before deadlines.